OpenStudy (anonymous):

If logb 3 = .23 and logb 8 =.71, find logb 24. Help

OpenStudy (anonymous):

\[\log (24) = \log (8 \times 3) = \log 8 + \log 3\]

OpenStudy (anonymous):

\[\log_b (24) = 0.23 + 0.71\]

OpenStudy (anonymous):

.94

OpenStudy (anonymous):

that's your answer

OpenStudy (anonymous):

are you familiar with how to condense logarithms?

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

do you have a question about it?

OpenStudy (anonymous):

yes i have been struggling with the following problem:

OpenStudy (anonymous):

if it's a new question you should post it as a new question

OpenStudy (anonymous):

and i also have to go now sorry

OpenStudy (anonymous):

thanks